# Euclid preparation

### v1

@@ 6 Validation
Figure 12: Posteriors of parameter inference obtained by fitting the Flagship (black and blue lines) and one of the EuclidLargeMocks (red and green lines) with cuts at $k<0.2$$h\,\text{Mpc}^{-1}$ (lower-left corner with red and black lines) and $k<0.3$$h\,\text{Mpc}^{-1}$ (upper-right corner with blue and green lines).
@@
The accuracy with which luminosity-dependent galaxy bias is reproduced is shown in Fig. 10. We report here the monopole of the power spectrum for flux cuts $f_{{\rm H}\alpha}>0.5f_{0}$, $f_{0}$, and $1.5f_{0}$, in the first and last redshift bin ($z\in[0.9,1.1]$ and $z\in[1.5,1.8]$). Luminosity-dependent bias is well reproduced, although at the lowest flux we notice an underestimation of the monopole that would require a more sophisticated implementation of CM to be removed; the number of galaxies detected in the actual survey at such low fluxes is expected to be small, so this discrepancy is unlikely to be a significant problem.
@@ 7 Prospects and conclusions
We presented the largest collection of simulated dark matter haloes in the light-cone ever produced, and a corresponding set of 3500$+$1000 galaxy mock catalogues of the Euclid spectroscopic sample that cover each an area of 2763 ${\rm deg}^{2}$, larger than the planned area of DR1 in Euclid Collaboration: Scaramella et al. (2022) and roughly one fifth of the final EWS. The simulations were produced with the Pinocchio code for the generation of approximate catalogues of dark matter haloes, in a numerical effort of $\sim$ 5 000 000 core-h that produced $\sim$ 300 TB of halo catalogues. The galaxy mock catalogues were obtained by extracting an HOD from the Flagship mock catalogue described in the Flagship paper, and calibrating it to absorb the little differences in the clustering and number density of Pinocchio and simulated haloes.
We validated these sets of galaxy mock catalogues by comparing number densities, power spectra and 2-point correlation functions, of haloes and galaxies with different halo mass or line flux cuts and in four redshift bins from $z=0.9$ to $1.8$, with the same measurements of the Flagship spectroscopic catalogue. We found that the spectroscopic sample drawn from the Flagship mock catalogue is consistent with being a realisation drawn from one of our sets, and this is testified by the reasonable values of $\chi^{2}_{\rm r}$ obtained using the Flagship multipoles as a data vector and the EuclidLargeMocks for the model and the covariance. We identified two main limitations of our mocks, in particular the Geppetto set is affected by the relatively small box size, replicated several times to fill the light-cone volume. This results in spurious oscillations of the galaxy power spectrum that can be mitigated by performing the measurements in units of the fundamental mode of the simulation box. This problem does not affect the EuclidLargeMocks, where the survey volume is contained in the simulation box. Also, For both sets we noticed an overestimation of the power spectrum quadrupole at $k>0.2$$h\,\text{Mpc}^{-1}$. Conversely, we noticed a flare of the quadrupole of the Flagship galaxies on the largest scale that is consistent with being a statistical fluctuations, while some blips in the Flagship galaxy number density are due to a known issue in the interpolation of internal extinction needed to compute the $\mathrm{H}\,\alpha$ line flux.
As a test of this consistency, we verified that using the power spectrum of a single EuclidLargeMocks or the Flagship mock for inferring parameters in a maximum likelihood fit, where the numerical covariance is obtained from the EuclidLargeMocks, gives consistent posteriors for all the cosmological and nuisance parameters. This remains true even when pushing the scale cut to $k<0.3$$h\,\text{Mpc}^{-1}$, showing that the overestimation of the quadrupole does not bias cosmological inference. This fully demonstrates the usability of these mock catalogues for analysing Euclid’s DR1.
The EuclidLargeMocks are being extensively used in many Euclid preparation papers, especially in the papers of the Observational Systematics Key Project presented by Euclid Collaboration: P. Monaco et al. (in prep.), and in the papers of the Organisation Unit for Level-3 products, in particular for testing the estimator of the power spectrum (Euclid Collaboration: Salvalaggio et al., in prep.). Simulated catalogues of dark matter haloes on the light-cone are going to me made available at https://adlibitum.oats.inaf.it/pierluigi.monaco/euclid_mocks.html by August 2026, Euclid mock catalogs will be released together with the Flagship mock catalogue.
@@ Appendix A Calibration of Pinocchio and Flagship halo masses
where $M_{\rm AM,12}:=\operatorname{log_{10}}(M_{\rm AM}/10^{12}\,h^{-1}\,M_{\odot})$ and $M_{\rm pin,12}:=\operatorname{log_{10}}(M_{\rm pin}/10^{12}\,h^{-1}\,M_{\odot})$ The coefficients are given in Table 2 for V4 and V5 of the code.
A brute-force determination of CM would imply to measure power spectra for hundreds of mock catalogues for many halo mass thresholds and redshifts; we resort to a more convenient procedure by assuming that the relation between AM and CM masses is well represented by a power law with redshift-dependent coefficients. This leaves us with two parameters, that are calibrated using the galaxy catalogues. We require that, at $k<0.2$$h\,\text{Mpc}^{-1}$, the average clustering amplitude of 100 Pinocchio galaxy catalogues with $f_{{\rm H}\alpha}>f_{0}$, measured in the four redshift bins with edges in $[0.9,1.1,1.3,1.5,1.8]$, is consistent with the Flagship one projected onto the same $30^{\circ}$ radius footprint, to within the predicted variance. On smaller scales, that are hardly used by the standard inference methods but are necessary to constrain the one-halo term, we require that the clustering amplitude is always reproduced to better than 10%. After a trial-and-error procedure based on 100 catalogues, we find that the CM masses of the Flagship mock is well reproduced with the present scaling

### v3

@@ 6 Validation
[…]
@@
The accuracy with which luminosity-dependent galaxy bias is reproduced is shown in Fig. 10. We report here the monopole of the power spectrum for flux cuts $f_{{\rm H}\alpha}>0.5f_{0}$, $f_{0}$, and $1.5f_{0}$, in the first and last redshift bin ($z\in[0.9,1.1]$ and $z\in[1.5,1.8]$). Luminosity-dependent bias is well reproduced, although at the lowest flux we notice an underestimation of the monopole that would require a more sophisticated implementation of CM to be removed (see the discussion in the Appendix); the number of galaxies detected in the actual survey at such low fluxes is expected to be small, so this discrepancy is unlikely to be a significant problem.
@@ 7 Prospects and conclusions
We presented the largest collection of simulated dark matter haloes in the light-cone ever produced, and a corresponding set of 3500++1000 galaxy mock catalogues of the Euclid spectroscopic sample that cover each an area of 2763 ${\rm deg}^{2}$, larger than the planned area of DR1 in Euclid Collaboration: Scaramella et al. (2022) and roughly one fifth of the final EWS. The simulations were produced with the Pinocchio code for the generation of approximate catalogues of dark matter haloes, in a numerical effort of $\sim$ 5 000 000 core-h that produced $\sim$ 300 TB of halo catalogues. The galaxy mock catalogues were obtained by extracting an HOD from the Flagship mock catalogue described in the Flagship paper, and calibrating it to absorb the little differences in the clustering and number density of Pinocchio and simulated haloes.
[…]
These mocks were produced for Euclid’s preparation. In particular, the EuclidLargeMocks are being extensively used in many Euclid preparation papers, especially in the papers of the Observational Systematics Key Project presented by Euclid Collaboration: P. Monaco et al. (in prep.), and in the papers of the Organisation Unit for Level-3 products, in particular for testing the estimator of the power spectrum (Euclid Collaboration: Salvalaggio et al., in prep.). Because they are calibrated on the Flagship simulation and not on real data, the scope of the galaxy mock catalogues is limited to such preparatory work. The halo light-cones will be reprocessed with an updated HOD to produce more realistic DR1-like catalogues, and will thus have a higher longevity. While nothing prevents us to use the EuclidLargeBox light-cones for the whole survey, for DR2 and DR3 we are planning an even more extended suite of simulations run with a code that improves on the scalability of the fragmentation part, and possibly with a larger simulation volume and better mass resolution.
8 Data availability
Simulated catalogues of dark matter haloes on the light-cone are available at https://adlibitum.oats.inaf.it/pierluigi.monaco/euclid_mocks.html, Euclid mock catalogues are available through the CosmoHub platform (Carretero et al. 2017; Tallada et al. 2020), https://cosmohub.pic.es.
@@ Appendix A Calibration of Pinocchio and Flagship halo masses
where $M_{\rm AM,12}:=\operatorname{log_{10}}(M_{\rm AM}/10^{12}\,h^{-1}\,M_{\odot})$ and $M_{\rm pin,12}:=\operatorname{log_{10}}(M_{\rm pin}/10^{12}\,h^{-1}\,M_{\odot})$ . The coefficients are given in Table 2 for V4 and V5 of the code.
[…]
@@
This CM guarantees that halos containing the bulk of $\mathrm{H}\,\alpha$ galaxies, namely those with $M_{\rm h}\sim 10^{12}$ $h^{-1}\,M_{\odot}$, have the same clustering level once the matching masses are used to define the mass cuts. However, as discussed in Munari et al. (2017b), halo bias is not recovered with sub-percent accuracy because halo construction is not perfect, and this weakens the clustering signal making its amplitude more similar to that of the matter field. This effect is stronger for smaller halos, while massive halos are reconstructed with higher accuracy. Figure 14 shows the halo power spectrum monopole in the first redshift bin for three mass cuts, where Pinocchio and Flagship mass cuts obey the CM relation. While halo clustering is very well recovered at $\sim$$10^{12}$ $h^{-1}\,M_{\odot}$, the correction is not strong enough for smaller halos while it is too strong at larger masses. This is in line with the trend shown in Fig. 10. A more sophisticated clustering matching procedure would easily correct for this difference, but this mismatch has a minor impact on the mock catalogues.
@@
Figure 14: Power spectrum monopole of halos, with $z\in[0.9,1.1]$, in the Flagship mock and in 50 EuclidLargeMocks for three distinct mass cuts that obey the CM relation.